OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation 1: m = 8 + 2n Equation 2: 3m = 4 + 4n Step 1: −3(m) = −3(8 + 2n) [Equation 1 is multiplied by −3.] 3m = 4 + 4n [Equation 2] Step 2: −3m = −24 − 6n [Equation 1 in Step 1 is simplified.] 3m = 4 + 4n [Equation 2] Step 3: −3m + 3m = −24 − 6n + 4n [Equations in Step 2 are added.] Step 4: 0 = −24 − 2n Step 5: n = −12 In which step did the student first make an error? A. Step 4 B. Step 3 C. Step 2 D. Step 1

OpenStudy (anonymous):

i think its step 4 because if you figure out the problem its 0 -2*-12=24 -24+24=0

OpenStudy (anonymous):

am i correct

OpenStudy (anonymous):

i have no clue cx